Please help me, I'm getting every solution but the correct one for this problem, thank you!
6x+7-4x=8+7x-26

Answer by nerdybill(7384)

6x+7-4x=8+7x-26
combine like-terms:
2x+7=7x-18
subtract 2x from both sides:
7=5x-18
add 18 to both sides:
25=5x
divide both sides by 5:
5 = x